# Natural inflation in Palatini $F(R)$

### v2

@@ 2 Palatini $F(R)$ theories
Finally, the number of e-folds can be expressed as follows:
@@ 3 Natural inflation for $n\leq 2$
which is positive and monotonically increasing for $\zeta>0$, provided that $1<n\leq 2$. The case $n=2$ was already studied explicitly in [32, 15]. Therefore, we focus on the range $1<n<2$. To obtain expressions for the CMB observables, we adopt the formalism derived in [38] and briefly outlined in Section 2. In terms of $\zeta_{N}$, we then have:
@@
Increasing the number of $e$-folds to $N_{e}=60$ (Fig. 3) shifts the predictions toward lower $r$ and slightly higher $n_{s}$, allowing the $M=6$ predictions to fall within the 95% confidence regions of the BICEP/Keck combination. Compatibility with the ACT combination at $\lesssim 2\sigma$ is achieved for $M\gtrsim 10$ and $n\gtrsim 7/4$, within an $n$-dependent range of $\alpha$, roughly between $10^{6}$ and $10^{8}$. Agreement with both combinations is realized for $n\gtrsim 7/4$, $10<M<500$, and $\alpha\sim 10^{7}$$10^{8}$.
@@ 5 Conclusions
We have analyzed natural inflation within the Palatini $F(R)$ framework, considering $F(R)=R+\alpha R^{n}$. Our results show that for $n<2$, the model can improve upon the predictions of standard natural inflation, particularly for $M\gtrsim 10$ and $7/4\lesssim n\leq 2$, with $\alpha$ lying within an appropriate $n$-dependent range. In this regime, the predictions for $r$ and $n_{s}$ can satisfy the constraints from the BICEP/Keck and/or ACT combinations.
For $n>2$, however, the allowed values of $\alpha$ are constrained by the observed scalar amplitude $A_{s}=2.1\cdot 10^{-9}$ and by the consistency of the equation of motion of the auxiliary field (Eq. 2.8), which limits the contribution of the higher-order term to the inflationary dynamics. Consequently, $r$ remains relatively high and $n_{s}$ does not shift sufficiently, preventing full agreement with observational constraints. Partial agreement can only be achieved for $n\rightarrow 2$, if $M\gtrsim 6$ and $\alpha\gtrsim 10^{8}$.
@@ Acknowledgments
This work was supported by the Estonian Research Council through grants PRG1055, RVTT3, and RVTT7, as well as the CoE program TK202 “Foundations of the Universe”. This article is based on work carried out within COST Actions COSMIC WISPers CA21106 and CosmoVerse CA21136, supported by COST (European Cooperation in Science and Technology).

### v3

@@ 2 Palatini $F(R)$ theories
Finally, the number of $e$-folds of remaining inflation can be expressed as follows:
@@ 3 Natural inflation for $n\leq 2$
which is positive and monotonically increasing for $\zeta>0$, provided that $1<n\leq 2$. We do not consider the case $n<1$ as we want that $F(R)\sim R$ as $R\rightarrow 0$, in order to recover General Relativity in the low energy limit. The case $n=2$ was already studied explicitly in [32, 15]. Therefore, we focus on the range $1<n<2$. To obtain expressions for the CMB observables, we adopt the formalism derived in [38] and briefly outlined in Section 2. In terms of $\zeta_{N}$, we then have:
@@
Increasing the number of $e$-folds to $N_{e}=60$ (Fig. 3) shifts the predictions toward lower $r$ and slightly higher $n_{s}$, allowing the $M=6$ predictions to fall within the 95% confidence regions of the BICEP/Keck combination. Compatibility with the ACT combination at $\lesssim 2\sigma$ is achieved for $M\gtrsim 10$ and $n\gtrsim 7/4$, within an $n$-dependent range of $\alpha$, roughly between $10^{6}$ and $10^{8}$. Agreement with both combinations is realized for $n\gtrsim 7/4$, $10<M<500$, and $\alpha\sim 10^{7}$$10^{8}$. Our analysis for $1<n\leq 2$ shows that inflationary scenarios consistent with current observations are obtained only for $n$ values quite close to $2$. The parameter $n$ controls how strongly the higher-order curvature term flattens the Einstein frame scalar potential $U$, and therefore suppresses tensor-to-scalar ratio $r$. As $n$ approaches $2$ from below (and $\alpha$ increases), the potential progressively acquires a plateau-like profile (see Fig. 4), similar to that found in $R^{2}$ inflation, and leads to predictions that are consistent with current CMB data.
@@ 5 Conclusions
We have analyzed natural inflation within the Palatini $F(R)$ framework, considering $F(R)=R+\alpha R^{n}$. Our results show that for $n<2$, the model can improve upon the predictions of standard natural inflation, particularly for $M\gtrsim 10$ and $7/4\lesssim n\leq 2$, with $\alpha$ lying within an appropriate $n$-dependent range. In this regime, the predictions for $r$ and $n_{s}$ can satisfy the constraints from the BICEP/Keck and/or ACT combinations. For $n>2$, however, the allowed values of $\alpha$ are constrained by the observed scalar amplitude $A_{s}=2.1\cdot 10^{-9}$ and by the consistency of the equation of motion of the auxiliary field (Eq. 2.8), which limits the contribution of the higher-order term to the inflationary dynamics. Consequently, $r$ remains relatively high and $n_{s}$ does not shift sufficiently, preventing full agreement with observational constraints. Partial agreement can only be achieved for $n\rightarrow 2$, if $M\gtrsim 6$ and $\alpha\gtrsim 10^{8}$.
Clearly, the $n<2$ is strongly favored over the $n>2$ case. However, the $n<2$ scenario still requires relatively tuned values for $n$ which are less natural choices than just $n=2$. On the other hand, with $7/4<n<2$, larger $n_{s}$ area can be covered at a given $M$, by varying $\alpha$.
An additional implication of our analysis is that a viable inflationary phase occurs only if the axion decay constant is transPlanckian. This requirement is a well known shortcoming of natural inflation scenarios, which cannot be solved by our construction. Realizing an effective super-Planckian scale ordinarily demands additional structure, such as alignment mechanisms, kinetic mixing effects, or multiple axion fields (see e.g. [48] and refs therein). Such a completion is beyond the purpose of our work and reserve a comprehensive ultraviolet embedding for future analyses.
@@ Acknowledgments
This work was supported by the Estonian Research Council through grants PRG1055, RVTT3, RVTT7, TARISTUS24-TK10, TARISTUS25-TK3 and the CoE program TK202 “Foundations of the Universe”. This article is based on work carried out within COST Actions COSMIC WISPers CA21106 and CosmoVerse CA21136, supported by COST (European Cooperation in Science and Technology).